Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

A board game has a spinner with equal sections, numbered 1 through 10. In order to win the game, Mae needs to spin at least a 6 on each of her last two turns. What's the probability of this happening?

OpenStudy (anonymous):

10 sections

OpenStudy (anonymous):

If it's numbered 1 thru 10, then the probability would be 1/10*1/10 = 1/100. That would be 1%

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

thanks

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!